OpenStudy (anonymous):

The ages of trees in a forest are normally distributed with a mean of 25 years and a standard deviation of 4. Approximately what percent of the trees are between 20 and 30 years old?

OpenStudy (anonymous):

z2 = (30-25)/4 = 5/4 = 1.25 z1 = (20-25)/4 = - 1.25 the fraction of the normal distribution between these two z scores is [from the cumulative normal distribution] 0.8944-.1056 = .7888 or 79%.